# Ab Initio Free Energy Surfaces for Coupled Ion-Electron Transfer

### v1

@@ Abstract
The Marcus theory of electron transfer assumes that diabatic energy gaps are sampled from a single ensemble. This assumption can break down in spatially anisotropic environments, such as electrochemical interfaces or biomolecular structures, where distinct solvent ensembles arise along a collective variable describing the anisotropy. Treating this collective variable as an additional reaction coordinate linearly independent from the Marcus reaction coordinate, we develop a formalism that enables calculation of the resulting Coupled Ion–Electron Transfer (CIET) free-energy surface directly from constrained ab initio trajectories. Applied to CO2 redox on a gold electrode, this method reveals strong coupling to the anisotropy, predicting significantly different activation barriers compared to either coordinate alone.
@@
Despite the emergence of Marcus theory in recent decades as a physically correct way to view ET reactions, the Butler–Volmer (BV) equation, which has empirically expressed Tafel’s law for current-voltage relationships of reaction-limited currents for nearly a century, assumes diabatic free energy curves that are linear instead of quadratic [14, 15]. In CIET, this discrepancy is resolved by postulating a 2D free energy surfaces which can in general have an arbitrary structure, but in the simplest approximation is linear when cross-sectioned along the ion-transfer CV and quadratic along the Marcus ET coordinate [7]. The result is that the CIET rate expression reduces to the Marcus–Hush–Chidsey (MHC) equation [16, 17, 18] in the limit where the reorganization energy dominates, and reduces to the BV equation in the limit where the barrier along the classical IT coordinate dominates [7]. The fact that viewing an electron transfer reaction as coupled to an anisotropic CV not only explains experimental reaction rates but also unifies physical equations correct in different regimes suggests the utility of such an approach, and suggests the possibility that a related framework can be generalized to symmetry-broken CVs other than IT at intercalation cathodes, finding broader applications across condensed phase physical chemistry [19, 20]. Practical analyses using these multidimensional surfaces may also become more feasible with concurrent improvements in sampling methods and machine-learned potentials [21, 22, 23, 24, 25, 26].
We note that ab initio computation of CIET parameters has recently been pioneered by Stenlid et al. [11] on a model lithium-cobalt-oxide battery interface. While remarkably effective in qualitatively reproducing experimental kinetic data, these calculations were conducted at 0 K, leaving entropic contributions to the free energy unaddressed. Furthermore, the study stopped short of computing the full 2D landscape, making assumptions, for example, that the reorganization energy was independent of the IT coordinate, and that adiabatic corrections to the landscape were insignificant.
@@
Here we present a formalism to compute finite-temperature 2D free-energy surfaces for ET coupled to a classical CV, expressed entirely in terms of quantities derived from constrained ab initio trajectories. We incorporate adiabatic corrections to the landscape as well, which prove to be significant [16]. We illustrate the method by applying it to CO2 redox on a gold electrode, and we find that the transition state (TS) of the 2D landscape can be significantly lower than when computed from the ET coordinate or CV alone. Although this demonstration assumes harmonic and symmetric ET, the framework allows this assumption to be readily relaxed, enabling generalization to diverse systems.
@@
Constructing the 2D Free Energy Surface. Within the Born–Oppenheimer approximation, an ab initio trajectory constrained to a diabatic electronic state can be modeled using constrained density functional theory (cDFT), as developed by Van Voorhis and co-workers [27, 28, 29] and implemented in various software packages [30, 31, 32]. Consider a simulation of atoms with the electronic state fixed to , where denotes the nuclear configuration. Let denote the nuclear trajectory under this constraint. At steady state, the probability of finding the system in configuration is given by the time-independent distribution . Letting denote the initial and final (i.e. oxidized and reduced) diabatic states of the electron transfer, the ET coordinate of is defined by the energy gap [4, 33, 34]
@@
Acknowledgements. This work was supported by the MIT Faculty Research Innovation Fund (FRIF) through the generosity of Irv and Melinda Simon. E.A. is grateful to Ezra Alexander, Mohammadhasan Dinpajooh, Beck Hanscam, Ziyue Jiang, Matthias Kick, Abraham Nitzan, Joakin Stenlid, Leah Weisburn, Junghyun Yoon, and Chao Zhang for useful discussions.

### v3

@@ Abstract
The Marcus theory of electron transfer assumes that diabatic energy gaps are sampled from a single ensemble. This assumption can break down in spatially anisotropic environments, such as Faradaic reactions at electrochemical interfaces, where distinct solvent ensembles arise along a collective variable describing the anisotropy. Treating this collective variable as an additional reaction coordinate linearly independent from the Marcus reaction coordinate, we develop a formalism that enables calculation of the resulting Coupled Ion–Electron Transfer (CIET) free-energy surface directly from constrained ab initio trajectories. Applied to CO2 redox on a gold electrode, this method reveals strong coupling to the anisotropy, predicting significantly different activation barriers compared to either coordinate alone.
@@
Despite the emergence of Marcus theory in recent decades as a physically correct way to view ET reactions, the Butler–Volmer (BV) equation, which has empirically expressed Tafel’s law for current-voltage relationships of reaction-limited currents for nearly a century, assumes diabatic free energy curves that are linear instead of quadratic [14, 15]. In CIET, this discrepancy is resolved by postulating a 2D free energy surfaces which can in general have an arbitrary structure, but in the simplest approximation is linear when cross-sectioned along the ion-transfer CV and quadratic along the Marcus ET coordinate [7]. The result is that the CIET rate expression reduces to the Marcus–Hush–Chidsey (MHC) equation [16, 17, 18] with an added IT prefactor in the limit where the reorganization energy dominates, and reduces to the BV equation in the limit where the barrier along the classical IT coordinate dominates [7]. The fact that viewing an electron transfer reaction as coupled to an anisotropic CV not only explains experimental reaction rates but also unifies physical equations correct in different regimes suggests the utility of such an approach, and suggests the possibility that a related framework can be generalized to symmetry-broken CVs other than IT at intercalation cathodes, finding broader applications across condensed phase physical chemistry [19, 20]. Practical analyses using these multidimensional surfaces may also become more feasible with concurrent improvements in sampling methods and machine-learned potentials [21, 22, 23, 24, 25, 26].
@@
We note that ab initio computation of CIET parameters has recently been explored by Stenlid et al. on a model lithium-cobalt-oxide battery interface [11]. While their reported parameters reproduced trends observed in experimental current-voltage relationships, we noticed several methodological inconsistencies in their study (see Supplementary Material for details). Here we present a physically consistent formalism to compute finite-temperature 2D free-energy surfaces for ET coupled to a classical CV, expressed entirely in terms of quantities derived from constrained ab initio trajectories. We incorporate adiabatic corrections to the landscape as well, which prove to be significant [16]. We illustrate the method by applying it to CO2 redox on a gold electrode, and we find that the transition state (TS) of the 2D landscape can be significantly lower than when computed from the ET coordinate or CV alone. Although this demonstration assumes harmonic and symmetric ET, the framework allows this assumption to be readily relaxed, enabling generalization to diverse systems.
@@
Constructing the 2D Free Energy Surface. Within the Born–Oppenheimer approximation, an ab initio trajectory constrained to a diabatic electronic state can be modeled using constrained density functional theory (cDFT), as developed by Van Voorhis and co-workers [27, 28, 29] and implemented in various software packages [30, 31, 32]. Consider a simulation of atoms with the electronic state fixed to , where denotes the nuclear configuration. Let denote a nuclear trajectory under this constraint. At steady state, the probability of finding the system in configuration is given by the time-independent distribution . Letting denote the initial and final (i.e. oxidized and reduced) diabatic states of the electron transfer, the ET coordinate of is defined by the energy gap [4, 33, 34]
@@
Acknowledgements. This work was supported by the MIT Faculty Research Innovation Fund (FRIF) through the generosity of Irv and Melinda Simon. E.A. is grateful to Ezra Alexander, Mohammadhasan Dinpajooh, Beck Hanscam, Ziyue Jiang, Matthias Kick, Abraham Nitzan, Joakim Stenlid, Leah Weisburn, Junghyun Yoon, and Chao Zhang for useful discussions.